# SurgSora: Decoupled RGBD-Flow Diffusion Model for Controllable Surgical Video Generation

### v1

@@ Abstract
Medical video generation has transformative potential for enhancing surgical understanding and pathology insights through precise and controllable visual representations. However, current models face limitations in controllability and authenticity. To bridge this gap, we propose SurgSora, a motion-controllable surgical video generation framework that uses a single input frame and user-controllable motion cues. SurgSora consists of three key modules: the Dual Semantic Injector (DSI), which extracts object-relevant RGB and depth features from the input frame and integrates them with segmentation cues to capture detailed spatial features of complex anatomical structures; the Decoupled Flow Mapper (DFM), which fuses optical flow with semantic-RGB-D features at multiple scales to enhance temporal understanding and object spatial dynamics; and the Trajectory Controller (TC), which allows users to specify motion directions and estimates sparse optical flow, guiding the video generation process. The fused features are used as conditions for a frozen Stable Diffusion model to produce realistic, temporally coherent surgical videos. Extensive evaluations demonstrate that SurgSora outperforms state-of-the-art methods in controllability and authenticity, showing its potential to advance surgical video generation for medical education, training, and research. See our project page for more results: surgsora.github.io.
@@
Endoscopic Video Diffusion Model Video Generation.
footnotetext: Project Lead;   Corresponding Author.
@@ 1 Introduction
Generative artificial intelligence (GAI) has achieved significant success in medical scenarios, including vision-language understanding [36], image restoration [6, 3], data augmentation [37], and medical report generation [20], advancing computer-aided diagnosis and intervention [2, 7]. Recently, researchers have explored video generation in endoscopic scenarios [23, 40, 18], where realistic dynamic videos offer high-quality resources to support clinician training, medical education, and AI model development. In particular, the controllability of video content—such as the motion of surgical instruments and tissues—becomes crucial for endoscopic surgical video generation [39]. Controllable generation enables dynamic and realistic surgical scenarios based on simple instructions, offering valuable applications in medical training. Furthermore, controllable video generation addresses the scarcity of annotated surgical data, reduces labeling costs, and enhances model generalization, accelerating downstream AI model development and deployment.
General scenarios of controllable video generation using diffusion models have been extensively explored [32], where various control signals—such as motion fields or flow-based deformation modules—are injected through specific parsers to produce videos with desired features and structures [27, 46]. While these approaches enable sophisticated editing of motion patterns, prior works on medical video generation have primarily focused on achieving visually plausible and temporally coherent outputs through effective spatiotemporal modelling [23, 40]. However, the crucial aspect of controllability—specifically for surgical videos—remains largely underexplored. Existing methods, such as Surgeon [8], rely on text descriptions to control video generation, but simple textual input often fails to capture the intricate and dynamic details of surgical procedures, limiting the precision of generated content.
To address this gap, we focus on controllable surgical video generation, where the primary challenge lies in accurately modeling the motion of surgical instruments and tissues based on intuitive user instructions. Given a single surgical image serving as the first frame, we allow users to specify motion directions through a straightforward process akin to direct clicking. This motion direction information is converted into sparse optical flow, which serves as a directive signal for the generation process. To facilitate controllable generation, we propose a novel framework that employs a dual-branch design to extract object-relevant RGB and depth features from the given first frame. These features are then warped using the optical flow data to represent the spatial information of the objects in subsequent frames. Leveraging our proposed multi-information guidance and decoupled flow mapper, our method effectively integrates targeted motion cues, detailed visual features, and object spatial dynamics, enabling the generation of realistic surgical videos with fine-grained motion and precise controllability. This approach not only fills the existing gap in controllable medical video generation but also opens new possibilities for high-fidelity, instruction-driven simulation of surgical scenarios. Our main contributions are summarized as follows.
  • We present the first work on motion-controllable surgical video generation using a diffusion model. This novel approach allows fine-grained control (both direction and magnitude) over the motion of surgical instruments and tissues, guided by intuitive motion cues provided by simple clicks.
  • We propose the Dual Semantic Injector (DSI), which integrates object-aware RGB-D semantic understanding. The DSI combines appearance (RGB) and depth information to better discriminate objects and capture complex anatomical structures, providing an accurate representation of the surgical scene.
  • We introduce the Decoupled Flow Mapper (DFM), which effectively fuses optical flow with semantic-RGB-D features at multiple scales. This fusion serves as the guidance conditions for a frozen Stable Video Diffusion model to generate realistic surgical video sequences.
  • We conduct extensive experiments on a public dataset, demonstrating the effectiveness of SurgSora in generating high-quality, motion-controllable surgical videos.
2 Related Works
2.1 Image-to-Video Generation
Researchers have explored generating videos from images and associated conditions, such as text descriptions [16] or motion control [1]. Controllability remains one of the most significant challenges in I2V generation research. A series of works explore incorporating multiple prompts (e.g., motion, clicks, text, and reference image) to provide more flexible control during video generation [43, 12, 26, 19]. MOFA-Video realizes controllable I2V with sparse motion hints (e.g., trajectories, facial landmarks) via domain-aware MOFA-Adapters to enable precise and diverse motion control across multiple domains [27]. Pix2Gif introduces explicit motion guidance, enabling users to define dynamic elements in the output, thus creating short, loopable animations [21]. Furthermore, ID-Animator [14] and I2V-Adapter [13] insert lightweight adapters into pretrained text-to-video models, employing cross-frame attention mechanisms to achieve efficient and effective I2V generation. In addition, several methods aim to further improve I2V generation performance and maintain high video quality and fidelity. For instance, PhysGen enhances the quality of generative models by using object dynamics and motion derived from physical properties as control conditions [25]. Meanwhile, ConsistI2V improves visual consistency in I2V generation by addressing temporal and spatial inconsistencies, ensuring high visual fidelity [28]. Although extensive research has been conducted on natural and animated scenes, the adaptation of these approaches to medical scenarios remains relatively unexplored and requires further investigation.
2.2 Medical Video Generation
Medical video generative models have been widely applied in various scenarios [24], such as ensuring privacy in echocardiogram videos [29], simulating disease progression [5], and editing the Ejection Fraction in ultrasound videos [30]. With advancements in diffusion model families, generalized text-to-video generative models have been explored for controllable generation in diverse medical contexts. For example, Bora, fine-tuned on custom biomedical text-video pairs, can respond to various medical-related text prompts [34]. In the field of endoscopy and surgery, Endora is an unconditional generative model designed as an endoscopy simulator, capable of replicating diverse endoscopic scenarios for educational purposes [23]. MedSora introduces an advanced video diffusion framework that integrates spatio-temporal Mamba modules, optical flow alignment, and a frequency-compensated video VAE. This framework enhances temporal coherence, reduces computational costs, and preserves critical details in medical videos [40]. Furthermore, Surgen generates realistic surgical videos based on text prompts [8], while Iliash et al. focused on generating videos with instrument-organ interaction with laparoscopic videos [18]. In our work, we aim to enable the model to generate realistic instrument motion in surgical scenes using simple motion cues (i.e., the direction of instrument movement).
@@ 3 Methodology
Figure 1: The pipeline of SurgSora: Segment features and depth images are generated from pre-trained models (SAM[22] and DAV2 [42]). a) denotes a Trajectory Controller (TC) module, decoding trajectories into sparse optical flow as the condition. b) illustrates the Dual Semantic Injector (DSI), which fuses RGB features and depth features with segment features and sends them into encoding blocks, respectively. c) Decoupled Flow Mapper (DFM) transforms images and depth features into optical flow separately to get decoupled flow features. The decoupled features are sent into the Multi-Scale Fusion Block for the following generation.
3.1 Overview
Our SurgSora framework, illustrated in Figure 1, comprises three key modules: the Dual Semantic Injector (DSI) introduced in Sec. 3.2, the Decoupled Flow Mapper (DFM) described in Sec. 3.3, and the Trajectory Controller (TC) module detailed in Sec. 3.4. Our model takes the first image frame $I_{RGB}\in\mathbb{R}^{3\times H\times W}$ as input. Based on $I_{RGB}$, the corresponding segmentation features $f_{seg}$ and depth image $I_{D}\in\mathbb{D}^{1\times H\times W}$ are generated from the pretrained Segment Anything Model [22] and the Depth Anything V2 [42]. The segment feature is injected into the RGB and Depth features in the DSI module to extract object-aware image features $f^{r}_{RGB}$ and depth features $f^{r}_{D}$ at multi-scales $r$. These features are then processed in the DFM module, where the optical flow $\theta\in\mathbb{O}^{(T-1)\times 2\times H\times W}$ (with $T$ as the total number of frames of the generated video), is resized and used to transform $f_{RGB}^{r}$ and $f_{D}^{r}$ independently. The transformed features are fused using the Multi-Scale Fusion (MSF) Block at different scales. These multi-scale fused features are then used as conditions for a frozen Stable Video Diffusion (SVD) model to generate the video.
@@ 3.2 Dual Semantic Injector
Traditional methodologies primarily rely on RGB images as input to create dynamic visual content. While effective in certain applications, this approach suffers from significant limitations in depth perception and scene understanding.Specifically, relying solely on RGB data complicates accurately capturing spatial relationships between objects, leading to deficiencies in visual coherence and object segmentation in generated videos. To address these challenges, we introduce the Dual Semantic Injector (DSI) module, a dual-branch architecture that enhances object awareness by integrating segmentation features into both the RGB and depth feature branches. Unlike traditional methods that depend solely on RGB images, we estimate and incorporate a depth map to provide crucial geometric cues. These cues improve the understanding of spatial relationships between objects and overall scene structure, making it especially beneficial for complex tasks like surgical video synthesis. Furthermore, to better discriminate between objects, object segmentation is leveraged to refine both RGB and depth features.
The segment features $f_{seg}$ are combined with RGB images $I_{RGB}$ and depth images $I_{D}$ by passing through two separate processors $\phi_{RGB}$ and $\phi_{D}$ for feature extraction and fusion, followed by two separate encoders for further encoding. The Dual Semantic Injector can be formulated as:
@@
Recall that the superscript $r$ indicates different scales of feature maps extracted by the encoders. This design uses a dual encoding method, which synchronizes and harmonizes the enhanced features from RGB and depth channels to optimize the overall representation. The injection of segmentation features enhances the semantic understanding compared with using the original RGB and depth features, significantly improving the discrimination of foreground and background, enhancing depth estimation, and ultimately contributing to more realistic and referenceable video predictions.
3.3 Decoupled Flow Mapper
Previous works [46, 27, 31] have demonstrated that the effectiveness of diffusion models can be significantly enhanced by adding additional information encoded into latent spaces. For these reasons, we employ a DFM module that bridges the spatial and sequential information of image and optical features to obtain spatial and temporal features for generating sequential videos. The object-aware RGB and depth features output by the DSI module are spatially transformed by the corresponding resized optical flow, respectively, elaborated as follows.
Let $f^{r}\in\mathbb{R}^{C_{r}\times H_{r}\times W_{r}}$ denote the output feature maps of the DSI module from either the RGB or the depth branch and $f^{r}(x,y)$ represent the feature at the position $(x,y)$. The optical flow $\theta\in\mathbb{O}^{(T-1)\times 2\times H\times W}$ is first resized to $\theta^{r}\in\mathbb{O^{\prime}}^{(T-1)\times 2\times H_{r}\times W_{r}}$ to match the size of $f^{r}$, and then used to spatially transform $f^{r}$ by applying the displacements $(\mathrm{d}x,\mathrm{d}y)$ provided in each frame $\theta^{r}_{t}$, in which $t\in[0,T)$ represents the current optical frame. The transformation is defined as:
$x^{\prime}=x+\mathrm{d}x,\qquad y^{\prime}=y+\mathrm{d}y.$ (2)
Here, $\mathrm{d}x$ and $\mathrm{d}y$ represent the displacements in the horizontal and vertical directions, respectively. Bilinear interpolation is used to estimate the updated pixel values at the new displacements $(\mathrm{d}x,\mathrm{d}y)$. The mapping procedure is given by:
$\hat{f}^{r}(x^{\prime},y^{\prime})=Interpolate(f^{r}(x,y)).$ (3)
Depth information typically captures geometry and spatial structure, while RGB information focuses on appearance and texture. To effectively leverage these complementary properties, we employ a decoupled-mapping method to independently spatially transform and extract frame features from the depth and RGB streams, and then integrate them via a Multi-Scale Fusion Block.
Multi-Scale Fusion Block (MSF) fuses the optical-flow-transformed RGB and depth features by concatenating them at different scales and then fusing them by two 3D convolution blocks and an activation block. The fusion process is expressed as:
@@ 3.3 Decoupled Flow Mapper
The fused feature $\check{f}_{fuse}^{r}$ is then used to assist a frozen Stable Video Diffusion Model for conditional video generation.
In sum, the integration of optical flow information enhances temporal continuity between frames, improving the smoothness and visual coherence of the generated videos. It enables the model to accurately capture scene dynamics and interactions, effectively interpreting complex motion. By fusing information from multiple modalities, the model achieves a more comprehensive understanding of scene depth and structure, maintaining visual authenticity while adapting to subtle changes. This multimodal strategy not only elevates video quality but also enriches detail and realism, ensuring a more accurate and dynamic representation of the scene.
@@ 3.4 Trajectory Controller
Figure 2: Workflow of Trajectory Controller (TC) module: First, the surgeon inputs the required surgical image as the first frame. Then, the surgeon clicks the image again to set the trajectories. The trajectories are sent into the Trajectories Decoder to get sparse optical for the following steps.
Surgical videos require more precision compared with natural videos, and generation with a single image makes it more difficult and non-referencable. Therefore, the Trajectory Controller (TC) is employed to enable custom trajectories as input for conditional motion generation. The pipeline of our TC module is shown in Figure 2, which employs a pre-trained trajectory decoder from [45]. The surgeon inputs the first frame image and then clicks to set trajectories. The trajectories and image will be encoded separately, concatenated together, and then decoded by the trajectory decoder into optical flows as a condition to guide the following generation. By involving the TC module, the quality of the generated video will be more referencable and convenient for generating customized surgical videos.
@@ 4 Experiment
4.1 Dataset and implementation details
We utilize the publicly available CoPESD dataset [38], which was collected from 20 videos using both conventional endoscopic submucosal dissection (ESD) and the DREAMS system [10], performed on in-vivo porcine models. The videos were recorded at a frame rate of 30 Hz with an original resolution of 1920 $\times$ 1080, which was cropped to 1300 $\times$ 1024. After an expert surgeon provided temporal annotations of ESD activities, video segments corresponding to submucosal dissection were separately extracted. We extract 21-frame video clips from the datasets, then resize them into $256\times 197$ resolution and pad them to $256\times 256$ resolution as the training and testing set. We employ the AdamW optimizer with a constant learning rate $2\times 10^{-5}$ for training. We set the batch size as 4 and trained the model on two RTX A6000 GPUs for 20000 iterations of training.
4.2 Comparison Methods and Evaluation Metrics
We compare performance against advanced video generation models on CoPESD datasets, including two unconditional generative video models, Endora [23] and StyleGAN-V [33]; and a conditional generative video model, MOFA-Video [27].
To evaluate the quality of the generated videos, we employ Fréchet Video Distance (FVD) [35] and Content-Debiased Fréchet Video Distance (CD-FVD) [11] to evaluate temporal consistency and the overall realism of the video, and Fréchet Inception Distance (FID) [15] and Inception Score (IS) [4] to evaluate the realism, fidelity, and diversity of individual frames. In addition, we also measure the frame consistency [9] by calculating the average CLIP cosine similarity of two consecutive frames. Meanwhile, for conditional video generation, we follow traditional pixel- or local-level video assessment metrics, including Peak Signal-to-Noise Ratio (PSNR) [17] and Structural Similarity Index (SSIM) [41], to assess the content and structural quality of the generated video against the ground truth. Furthermore, we include optical flow evaluation metrics, specifically the F1-epe and F1-all scores, to assess the consistency between the optical flow of the generated video and the given optical flow.
4.3 Results
4.3.1 Video Generation Evaluation and Visualization
Figure 3: Comparison of unconditional generation results trained on CoPESD dataset: Both Endora [23] and StyleGAN-V [33] models generated videos with different levels of noticeable gaps compared with the real surgical scene (tissues and instruments).
Table 1: Quantitative Comparisons on CoPESD Dataset [38]. The comparison contains both video consistency and video content qualities.

### v2

# SurgSora: Object-Aware Diffusion Model for Controllable Surgical Video Generation

@@ Abstract
Surgical video generation can enhance medical education and research, but existing methods lack fine-grained motion control and realism. We introduce SurgSora, a framework that generates high-fidelity, motion-controllable surgical videos from a single input frame and user-specified motion cues. Unlike prior approaches that treat objects indiscriminately or rely on ground-truth segmentation masks, SurgSora leverages self-predicted object features and depth information to refine RGB appearance and optical flow for precise video synthesis. It consists of three key modules: (1) the Dual Semantic Injector, which extracts object-specific RGB-D features and segmentation cues to enhance spatial representations; (2) the Decoupled Flow Mapper, which fuses multi-scale optical flow with semantic features for realistic motion dynamics; and (3) the Trajectory Controller, which estimates sparse optical flow and enables user-guided object movement. By conditioning these enriched features within the Stable Video Diffusion, SurgSora achieves state-of-the-art visual authenticity and controllability in advancing surgical video synthesis, as demonstrated by extensive quantitative and qualitative comparisons. Our human evaluation in collaboration with expert surgeons further demonstrates the high realism of SurgSora-generated videos, highlighting the potential of our method for surgical training and education. Our project is available at surgsora.github.io.
@@
Surgical Video Diffusion Model Video Generation.
footnotetext: * Equal Contribution;   Project Lead;   Corresponding Author.
@@ 1 Introduction
Surgical video generation has the potential to enhance medical education, clinician training, and AI-driven surgical analysis by providing realistic and controllable visual representations of complex procedures [2, 19, 15]. However, existing methods face two key challenges: visual authenticity and generation controllability. Different strategies have been proposed to generate high-quality endoscopic videos. Endora [9] is an endoscopy simulator capable of replicating diverse endoscopic scenarios while lacking precise controlling abilities. MedSora [24] focuses on past temporal coherence to achieve high-fidelity forward video generation, while predicting based on the optical flow of past frames is challenging due to the complexity of surgical scenaris. Moreover, most current models indiscriminately process entire scenes, failing to differentiate individual objects, leading to blurred boundaries and unrealistic motion dynamics. Iliash et al. [7] attempt to improve object awareness using segmentation masks, while this approach requires additional object-level annotations and hard object boundaries, limiting the ability to model smooth transitions and fine anatomical details.
Beyond realism, precise motion control is critical for replicating surgical actions such as tool manipulation and tissue interaction [22]. However, most existing methods lack this capability, making it difficult to generate videos that follow user-specified actions. In the context of general video generation, text prompts are a commonly used control method [10, 23, 2, 19, 7]. However, surgical scenarios often involve precise and detailed actions, which are difficult to fully capture through textual descriptions. This leads to a domain gap between language and motion, making it challenging for text-based guidance to express fine-grained surgical actions. In addition, surgical triplets (tissue-instrument-action) have also been used for video generation tasks [13, 14, 27], while triplets do not include specific motion and directional information, making it difficult to comprehensively represent the rich information contained in surgical videos.
To this end, we propose SurgSora, an object-aware trajectory-controlled RGB-Depth Flow Diffusion model that introduces self-predicted object-relevant RGB and depth features for precise object localization — eliminating the need for segmentation masks. Instead of relying on text-based control, SurgSora utilizes motion trajectories, offering fine-grained, user-directed control over object movements, enabling the generation of realistic and clinically relevant surgical videos. We can freely generate realistic videos by providing motion guidance using our SurgSora, offering abundant surgical training and education materials.
Our contribution can be summarized as: (i) We present the first work on motion-controllable surgical video generation using a diffusion model, which allows fine-grained control (both direction and magnitude) over the motion of surgical instruments and tissues, guided by intuitive motion cues provided by simple clicks. (ii) We propose the Dual Semantic Injector (DSI), which integrates object-aware RGB-D semantic understanding. The DSI combines appearance (RGB) and depth information to better discriminate objects and capture complex anatomical structures, providing an accurate representation of the surgical scene. (iii) We introduce the Decoupled Flow Mapper (DFM), which effectively fuses optical flow with semantic-RGB-D features at multiple scales. This fusion serves as the guidance conditions for a frozen Stable Video Diffusion model to generate realistic surgical video sequences. (iv) Extensive experiments on the CoPESD dataset demonstrate the effectiveness of SurgSora in generating high-quality, motion-controllable surgical videos. We further involve surgeons verifying the authenticity of our videos to justify their usefulness in medical use.
@@ 2 Methodology
Figure 1: SurgSora Pipeline: a) Trajectory Controller module encodes trajectories into sparse optical flow. b) Dual Semantic Injector merges RGB and depth features with segment features separately. c) Decoupled Flow Mapper maps RGB and depth features into optical flow, then sends to Multi-Scale Fusion Block as condition.
Our SurgSora framework, as illustrated in Fig. 1, comprises three key modules: the Dual Semantic Injector (DSI) introduced in Sec. 2.1, the Decoupled Flow Mapper (DFM) described in Sec. 2.2, and the Trajectory Controller (TC) module detailed in Sec. 2.3. Our model takes the first image frame $I_{RGB}\in\mathbb{R}^{3\times H\times W}$ as input. Based on $I_{RGB}$, the corresponding segmentation features $f_{seg}$ and depth image $I_{D}\in\mathbb{D}^{1\times H\times W}$ are generated from the pretrained Segment Anything Model [8] and the Depth Anything V2 [26]. The segment feature is injected into the RGB and Depth features in the DSI module to extract object-aware image features $f^{r}_{RGB}$ and depth features $f^{r}_{D}$ at multi-scales $r$. These features are then processed in the DFM module, where the optical flow $\theta\in\mathbb{O}^{(T-1)\times 2\times H\times W}$ (with $T$ as the total number of frames of the generated video), is resized and used to transform $f_{RGB}^{r}$ and $f_{D}^{r}$ independently. The transformed features are fused using the Multi-Scale Fusion (MSF) Block at different scales. These multi-scale fused features are then used as conditions for a frozen Stable Video Diffusion (SVD) model to generate the video.
@@ 2.1 Dual Semantic Injector
Traditional methodologies primarily rely on RGB images as input to create dynamic visual content. While effective in certain applications, this approach suffers from significant limitations in depth perception and scene understanding. Specifically, relying solely on RGB data complicates accurately capturing spatial relationships between objects, leading to deficiencies in visual coherence and object segmentation in generated videos. To address these challenges, we introduce the Dual Semantic Injector (DSI) module, a dual-branch architecture that enhances object awareness by integrating segmentation features into both the RGB and depth feature branches. Unlike traditional methods that depend solely on RGB images, we estimate and incorporate a depth map to provide crucial geometric cues. These cues improve the understanding of spatial relationships between objects and overall scene structure, making it especially beneficial for complex tasks like surgical video synthesis. Furthermore, to better discriminate between objects, object segmentation is leveraged to refine both RGB and depth features. The segment features $f_{seg}$ are combined with RGB images $I_{RGB}$ and depth images $I_{D}$ by passing through two separate processors $\phi_{RGB}$ and $\phi_{D}$ for feature extraction and fusion, followed by two separate encoders for further encoding. The Dual Semantic Injector can be formulated as:
@@
Recall that the superscript $r$ indicates different scales of feature maps extracted by the encoders. This dual encoding method synchronizes and harmonizes the enhanced features from RGB and depth channels to optimize the overall representation. The segmentation features enhance the semantic understanding compared with using the original RGB and depth features, significantly improving the discrimination of foreground and background, enhancing depth estimation, and ultimately contributing to more realistic and referenceable video predictions.
2.2 RGB-Depth Frame Mapper
Recent studies [30, 12, 17] have highlighted the advantages of integrating additional latent space information into diffusion models to enhance their output quality. Building on this insight, our Decoupled Flow Mapper (DFM) module integrates spatial and temporal data from image and optical features to generate sequential videos effectively. The DFM uses object-aware RGB and depth features from the Dual Semantic Injector (DSI) module, transformed spatially by resized optical flow to produce enriched video sequences.
Specifically, feature maps $f^{r}\in\mathbb{R}^{C_{r}\times H_{r}\times W_{r}}$ from the DSI module are modified according to the resized optical flow $\theta^{r}\in\mathbb{O^{\prime}}^{(T-1)\times 2\times H_{r}\times W_{r}}$, aligning them spatially for each frame. The features are shifted by $\mathrm{d}x$ and $\mathrm{d}y$, the displacements provided by $\theta^{r}_{t}$, with new positions computed as $x^{\prime}=x+\mathrm{d}x$ and $y^{\prime}=y+\mathrm{d}y$. Interpolation determines the new pixel values at these coordinates, effectively merging the depth information’s structural insights with the RGB data’s textural details. This decoupled approach ensures that each feature type is optimally utilized, enhancing the video generation process. The fused features from separate streams are then passed the Multi-Scale Fusion Block (MSF) for final video output. The MSF block shall fuse the optical-flow-transformed RGB and depth features by concatenating them at different scales and then fusing them with 3D convolution blocks and an activation block, which can be formulated as:
@@ 2.2 RGB-Depth Frame Mapper
The fused feature $\check{f}_{fuse}^{r}$ is then used to assist a pretrained Stable Video Diffusion (SVD) Model for conditional video generation. The integration of optical flow information substantially enhances the temporal continuity and smoothness for generation, enabling accurate capture of scenes and complex objective motions. By utilizing fused data, the model gains a richer understanding of scene depth and structure, ensuring visual authenticity and adaptability to nuanced changes. This approach improves video quality through diversity-aware fusion, providing a dynamic and precise scene representation.
@@ 2.3 Trajectory Controller
Surgical videos demand higher precision compared to natural scene videos, particularly because generating them from a single image introduces significant challenges and reduces referential accuracy. Due to this we employ trajectory for precision control, which employs a pre-trained trajectory decoder from [29]. The surgeon inputs the first frame image and then clicks to set trajectories. The trajectories and image will be encoded separately, concatenated together, and then decoded by the trajectory decoder into optical flows as a condition to guide the following generation. By involving the TC module, generated videos will be more referencable and convenient for generating customized surgical videos.
@@ 3 Experiment
Dataset and Implementation Details. We utilize the publicly available CoPESD dataset [21], which was collected from 20 videos using both conventional endoscopic submucosal dissection (ESD) and the DREAMS system [3], performed on in-vivo porcine models. The videos were recorded at a frame rate of 30 Hz with an original resolution of 1920 $\times$ 1080, which was cropped to 1300 $\times$ 1024. After an expert surgeon provided temporal annotations of ESD activities, video segments corresponding to submucosal dissection were separately extracted. We extract 21-frame video clips from the datasets, then resize them into $256\times 197$ resolution and pad them to $256\times 256$ resolution as the training and testing set. for training. For the training process, we use the AdamW [11] with a learning rate of $2\times 10^{-5}$, a batch size of 4, and train on two A6000 GPUs.
Comparison Methods and Evaluation Metrics. We assess the performance of our model against leading video generation models: the unconditional models Endora [9] and StyleGAN-V [18]; and the conditional model MOFA-Video [12]. For video quality evaluation, we employ metrics such as Fréchet Video Distance (FVD)[20] and Content-Debiased Fréchet Video Distance (CD-FVD)[4] to assess temporal consistency and video realism. Frame realism and diversity are evaluated using Fréchet Inception Distance (FID)[5] and Inception Score (IS)[1]. Frame consistency is measured by CLIP cosine similarity [16] between consecutive frames. For conditional generation, we employ PSNR [6] and SSIM [25] for content and structural accuracy, alongside optical flow metrics F1-epe and F1-all, to verify flow consistency with the input trajectories.
3.1 Experimental Results
Table 1: Quantitative Comparisons on CoPESD Dataset [21].
Models Discriminate Avg Acc Frame Consistency $\uparrow$ Video Optical-Flow